OpenStudy (anonymous):

Help with a true or false! O: For a trig function, y=f(x), then x=F^-1(y), I put false but I need an explanation on why that is.

OpenStudy (anonymous):

lets pick an example \[\sin(4\pi)=0\] but \[\sin^{-1}(0)\neq 4\pi\]

OpenStudy (anonymous):

it is because sine is is not a one to one function, so will not have an inverse unless you restrict the domain (hence the range of the inverse function)

OpenStudy (anonymous):

Thanks! I have a couple more that are similar, if you dont mind helping me with those too? @satellite73